# Intercept Cancer: Cancer Pre-Screening with Large Scale Healthcare Foundation Models

### v1

@@ Abstract
Cancer screening, leading to early detection, saves lives. Unfortunately, existing screening techniques require expensive and intrusive medical procedures, not globally available, resulting in too many lost would-be-saved lives. We present CATCH-FM, CATch Cancer early with Healthcare Foundation Models, a cancer pre-screening methodology that identifies high-risk patients for further screening solely based on their historical medical records. With millions of electronic healthcare records (EHR), we establish the scaling law of EHR foundation models pretrained on medical code sequences, pretrain compute-optimal foundation models of up to 2.4 billion parameters, and finetune them on clinician-curated cancer risk prediction cohorts. In our retrospective evaluation comprising of thirty thousand patients, CATCH-FM achieved strong efficacy (60% sensitivity) with low risk (99% specificity and Negative Predictive Value), outperforming feature-based tree models as well as general and medical large language models by large margins. Despite significant demographic, healthcare system, and EHR coding differences, CATCH-FM achieves state-of-the-art pancreatic cancer risk prediction on the EHRSHOT few-shot leaderboard, outperforming EHR foundation models pretrained using on-site patient data. Our analysis demonstrates the robustness of CATCH-FM in various patient distributions, the benefits of operating in the ICD code space, and its ability to capture non-trivial cancer risk factors. Our code will be open-sourced.
@@ 1 Introduction
Early cancer detection by cancer screening is one of the most effective ways to combat cancer [71]. Cancers detected at an early stage are treated with significantly improved patient outcomes [32; 46]. Recent medical advancements also significantly improved the curative rates for cancers detected at early stages [42; 61; 74; 85]. Routine screening with follow-up monitoring of high cancer risk patients is standard practice, enabling timely intervention and effective treatments [2; 31; 68].
@@
Despite its benefits, cancer screening remains underutilized [105], especially in populations with limited healthcare resources [99] due to reliance on invasive, resource-intensive procedures like medical imaging [67; 96]. Screening is more common for cancers with clear risk factors, like breast and colorectal [72], while cancers without early symptoms, such as pancreatic cancer, often progress silently and are detected late, with survival durations under one year [5].
@@
We present CATCH-FM: CATch Cancer early with Healthcare Foundation Models, a cancer pre-screening tool that identifies high-risk patients using only their medical history. CATCH-FM is pretrained on large-scale longitudinal EHR data and finetuned on clinician-curated cancer cohorts. It directly operates on precise medical codes (ICD), learning general medical patterns through next-code prediction and finetuned to capture cancer risk signals reflected in patient medical history [49; 50; 63]. Once trained, CATCH-FM can be deployed in EHR systems to predict cancer risk at low cost, supporting healthcare providers in deciding when and whom to screen.
@@
To facilitate the study of EHR foundation models in cancer pre-screening, we build NHIRD-Cancer, a cancer risk prediction benchmark, by sampling more than three million patients from the Taiwanese National Health Insurance Research Database (NHIRD) [35], a government de-identified and research-accessible EHR database. We allocate three million patients for pretraining, consisting of billions of medical events spanning two decades, and the remainder for finetuning and evaluation. We focus on three cancers based on their critical needs for early detection: pancreatic, liver, and lung cancers [46; 48; 85], and curate a clinically reliable training and testing set by matching cancer patients with control-to-case group patient cohorts.
@@
With billions of medical events available, we examine how compute budget (FLOPs), model size, and pretraining tokens affect performance. Our findings establish an EHR foundation model scaling law and confirm the benefit of large-scale pretraining on EHR data. Accordingly, we pretrain compute-optimal CATCH-FM models up to 2.4b parameters. When finetuned and evaluated for cancer risk prediction on NHIRD-Cancer, CATCH-FM consistently outperforms feature-based tree models and language models trained on the same data. Without loss of flexibility, it achieves strong high-risk prediction efficacy (60%+ sensitivity) with reliable classification of low-risk patients (99% specificity and Negative Predictive Value), often considered a minimum standard for screening techniques [7; 28]).
@@
On the pancreatic cancer risk prediction task from Stanford Medicine EHRSHOT benchmark [97], CATCH-FM outperforms the prior state-of-the-art EHR foundation model, CLMBR [97], pretrained on millions of on-site patient records, while CATCH-FM is trained on data from drastically different populations, disease prevalence, and coding systems (ICD vs. SNOMED). Our analyses further demonstrate the robustness of CATCH-FM across different patient cohorts, preexisting conditions, and pre-screening configurations. Our interpretability analyses [27] reveal that CATCH-FM identified not only known cancer risk factors but also non-trivial markers discovered in recent medical research.
@@ 2 Related Work
Cancer screening significantly improves prognosis and patient outcomes [46; 48; 85]. Advancements in cancer treatment have made cancers more treatable and potentially curable, if they are detected in early stages [16; 36; 46; 55]. Recent advances in AI-powered medical imaging, like CT scans with multimodal models, have enhanced cancer screening accuracy, sometimes exceeding human-level sensitivity [7; 11; 94; 98]. The challenge is that medical imaging is resource-intensive, inaccessible for under-served populations, and not widely performable [23; 87; 90; 91].
@@
Using electronic health records (EHR) to assess cancer risk is a promising path to improve cancer screening effectiveness and efficiency, i.e., by identifying patients with high cancer risk for healthcare professionals to make informed cancer screening decisions [49; 50]. Recent approaches have made attempts using feature-based machine learning models on large, task-specific datasets to detect cancer risks, and have shown the possibility of capturing cancer risk signals in EHR [62; 64].
@@
Large language models pretrained on medical corpora are effective on text-oriented medical tasks [17] such as medical question answering [73], clinical document summarization [57], and radiology report generation [80]. Recent research has also pretrained foundation models on structured EHR data [15; 26; 103], and explored their utilities in various clinical tasks [70; 79; 102]. A challenge for EHR foundation models is that many EHR datasets only cover a slice of patients’ healthcare records [24; 41; 66], e.g., MIMIC-IV mainly focuses on emergency departments and ICU encounters [41]. Foundation models pretrained on EHR slices are more effective in corresponding prediction tasks such as emergency department triage [79] and ICU readmission risk [40].
@@
Many tasks, like cancer risk prediction, require longitudinal EHR data to capture full patient history. Foresight uses two decades of data to pretrain Transformers for biomedical forecasting [47]. MOTOR trains a 143M-parameter model on time-to-event tasks using 2.7M patient records from 2014–2022, effectively predicting diagnosis time [75]. CLMBR, a 141M-parameter Transformer, is pretrained on 2.57M Stanford Medicine records over three decades [97], with 6.7K records released in the EHRSHOT benchmark for 15 few-shot prediction tasks [97].
@@
Due to privacy and ethical constraints, large-scale longitudinal EHR data are hard to release publicly. Most EHR foundation models are pretrained on one or two hospital sites and are limited in scale to around 100M parameters [75; 97]. Their advantage over feature-based models in cancer risk prediction remains unclear, for instance, while CLMBR performs well on procedure outcomes and lab predictions, it fails to outperform feature-based models in pancreatic cancer prediction on EHRSHOT [97].
@@ 3.1 NHIRD Preliminary
Overview. The National Health Insurance Research Database (NHIRD) include electronic health record (EHR) of over 99.99% of Taiwan population. It includes decades of de-identified records from all healthcare encounters under the National Health Insurance program, with diagnoses, prescriptions, and procedures [35]. Coding standards are applied uniformly across providers, ensuring consistency and correctness. A sample patient record is shown in Appendix Figure 6(b).
Availability. The sensitivity of longitudinal EHR data makes large-scale public access challenging [29]. Among existing longitudinal EHRs [47; 75; 97], NHIRD is among the most accessible for open research. Eligible institutions in Taiwan and their international collaborators can access NHIRD with IRB approval or exemption under standardized protocols. Access is regulated by the Ministry of Health and Welfare to ensure strong privacy protections [35; 53; 82]. See Appendices A and B for details on de-identification, data quality, and access to NHIRD.
@@
| Total | 315,827 | 294,516 | 310,205 |
| Training size (80%) | 252,662 | 235,613 | 248,164 |
| Validation size (10%) | 31,583 | 29,452 | 31,020 |
| Test size (10%) | 31,582 | 29,451 | 31,021 |
| Positive | 12,015 | 10,991 | 8,675 |
| Negative | 303,812 | 283,525 | 301,530 |
| Avg. history | 12.3 years | 13.8 years | 14 years |
@@ 3.2 NHIRD-Cancer Benchmark Curation
Data Statistics. We accessed a subset of NHIRD via collaboration with a Taiwanese medical school, with IRB exemption, as the study uses anonymized, non-human-subjects data. The dataset includes 3+ million patients from 1996–2013, covering 1.4 billion visits with diagnoses, treatments, and medications coded in ICD-9 [8]. It provides 15+ years of physician-confirmed, government-validated history and over 8 billion medical codes. Summary statistics are in Table 1, with preprocessing and additional details in Appendix A. Our subset aligns with full NHIRD averages (18 vs. 16.5 visits per patient over 15 years) and is comparable in scale to the 3.67M-patient records from Stanford medicine used in CLMBR [97], which is not publicly available.
Targeted Cancers. We focus on pancreatic, liver, and lung cancers, which have high mortality rates and increased risk of secondary cancer [21; 30; 39] where early detection offers substantial benefit [48; 85; 5; 32]. The screening task is a binary classification: predicting whether a patient will later develop the target cancer. As suggested by clinicians, we focus on predicting cancer one year after a clinical visit, enabling early intervention for proactive care. We set the available medical history per patient to be five years, following clinical standards.
Case and Control Group. Guided by clinicians, we rigorously follow the standard case-control study methodology and previous cancer screening studies on NHIRD [63; 49; 50] to create case-control dataset. We use the first occurance of target cancer type as the positive case, but note that patients may have other types of cancers, which we consider as the reoccurred group. The other patients are first cancer patients. Screening for both is evaluated. The (negative) control group includes patients with no cancer diagnoses, ensuring a clear distinction between cases and controls, and are matched with the case group as follows.
@@
Our case-control matching goes beyond standard demographics by incorporating factors such as history-length and timing. This approach ensures closer alignment between case and control patients, accounting for temporal and historical factors, ensuring the observed differences reflect the presence or absence of cancer, thereby enhancing reliability and validity.
Dataset Processing. We use three-digit ICD-9 codes 157, 155, and 162 to define pancreatic, liver, and lung cancers, respectively. Control patients must not have cancer-related codes from 140–239 (top level of Neoplasms). To support broader cancer types and maximize early detection [13], we use top-level ICD codes to capture more cases. Table 2 lists benchmark statistics. All cancer diagnoses are confirmed by licensed physicians and validated by the government. More details in Appendix A.
Availability. NHIRD-Cancer follows the same access control as NHIRD. Approved NHIRD users can request a copy through a lightweight approval process. All data processing and benchmark curation code will be open-sourced for use with other EHR datasets.
@@ 4.1 Model
Patient Representation. In longitudinal structured EHR databases, the patient records are often represented by a coding system, for example, ICD [8] or SNOMED [19], using a unique ID to encode each specific piece of information. A typical patient record can start with their demographic information, such as age and gender , followed by a sequence of chronologically ordered visits, . Each visit consists of medical events, , covering diagnoses, medications, procedures, and other clinical events.
@@
Architecture. CATCH-FM uses the standard decoder-only transformer architecture on top of the patient’s EHR token sequence . We use rotary positional embeddings (RoPE) [78] to encode positional information for visits:
@@ 5 Experimental Methodologies
Dataset. We use the NHIRD-Cancer benchmark described in Sec. 3.2 for the pretraining, finetuning, and in-domain evaluation. There is no patient overlap among the three. To evaluate model generalization, we use the EHRSHOT benchmark [97], containing 6,739 Stanford Medicine patients (1990–2023), primarily coded in SNOMED, CPT, and RxNorm. It introduces significant distribution shifts from NHIRD: different healthcare systems, different medical codings, and different populations with different cancer prevalence. We use EHRSHOT’s official pancreatic cancer risk prediction task following their few-shot setting (1–128 positives). It is the only available cancer risk prediction task.
As CATCH-FM is pretrained on NHIRD using ICD-9 and NHI codes, we map SNOMED to ICD-9 using official CMS and NLM mappings. While 38% of codes have one-to-one mappings (mostly SNOMED to ICD-9), the remaining 62% (e.g., RxNorm, CPT) are aligned to NHI codes via semantic matching on text descriptions [56; 81; 104]. Due to differences in drug, procedure, and medical codes, the average cosine similarity between EHRSHOT and top-1 matched NHIRD codes is 84.3%, posing a significant generalization challenge. Appendix C details the mapping process.
@@
Evaluation metrics. NHIRD-Cancer evaluations use AUPRC, specificity, sensitivity, and Negative Predictive Value (NPV). AUPRC captures the precision-recall trade-off in imbalanced tasks like cancer risk. Specificity, the ratio of true negatives to all negatives, reflects reliability in identifying low-risk patients. NPV measures confidence in negative results and is often a key metric for screening [28]. Sensitivity, the ratio of true positives to total positives, reflects how well the model flags cancer-risk patients. When possible, we evaluate sensitivity at 99% specificity, a common (pre)screening threshold [7], and a clinician-defined standard to reduce false positives, ensuring clinical trust. EHRSHOT uses AUPRC and AUROC as official metrics, and we follow exact setups.
@@
Baselines. We include standard tree-based models with bag-of-words as input [3]: XGBoost [12] and LightGBM [45], and well-established EHR Deep learning models [92]: StageNet [26] and RETAIN [15]. We compare with pretrained language models, by converting medical codes into text and finetuning BioGPT [58] and long input sequence Qwen2.5 [100] on the same data. For EHRSHOT, we directly compare CATCH-FM against methods reported in their leaderboard [97].
@@
Implementation Details. We implement CATCH-FM using a standard architecture [4], a decoder-only transformer with rotary embeddings and flash attention, and pretrain from scratch on NHIRD. Medical codes, demographics, and special tokens form a 185,138 token vocabulary, with sequences capped at 2,048 tokens. Longer histories are chunked without overlap during pretraining and truncated for fine-tuning. More architecture, hyperparameters, and training details are provided in Appendix E.

### v2

@@ Abstract
Cancer screening, leading to early detection, saves lives. Unfortunately, existing screening techniques require expensive and intrusive medical procedures, not globally available, resulting in too many lost would-be-saved lives. We present CATCH-FM, CATch Cancer early with Healthcare Foundation Models, a cancer pre-screening methodology that identifies high-risk patients for further screening solely based on their historical medical records. With millions of electronic healthcare records (EHR), we establish the scaling law of EHR foundation models pretrained on medical code sequences, pretrain compute-optimal foundation models of up to 2.4 billion parameters, and finetune them on clinician-curated cancer risk prediction cohorts. In our retrospective evaluation comprising of thirty thousand patients, CATCH-FM achieves strong efficacy, with 50% sensitivity in predicting first cancer risks at 99% specificity cutoff, and outperforming feature-based tree models and both general and medical LLMs by up to 20% AUPRC. Despite significant demographic, healthcare system, and EHR coding differences, CATCH-FM achieves state-of-the-art pancreatic cancer risk prediction on the EHRSHOT few-shot leaderboard, outperforming EHR foundation models pretrained using on-site patient data. Our analysis demonstrates the robustness of CATCH-FM in various patient distributions, the benefits of operating in the ICD code space, and its ability to capture non-trivial cancer risk factors. Our code will be open-sourced.
@@ 1 Introduction
Early cancer detection by cancer screening is one of the most effective ways to combat cancer (Siegel et al., 2024). Cancers detected at an early stage are treated with significantly improved patient outcomes Haue et al. (2024); Kim et al. (2024). Recent medical advancements also significantly improved the curative rates for cancers detected at early stages (Ju et al., 2024; Maru & Jaffee, 2024; Springfeld et al., 2023; Thiele et al., 2024). Routine screening with follow-up monitoring of high cancer risk patients is standard practice, enabling timely intervention and effective treatments (Altmayer et al., 2024; Gyawali & Booth, 2024; Rubinstein et al., 2024).
@@
Despite its benefits, cancer screening remains underutilized Zhang et al. (2022), especially in populations with limited healthcare resources Xu et al. (2024b) due to reliance on invasive, resource-intensive procedures like medical imaging (Rohatgi et al., 2020; Washington & Deville, 2020). Screening is more common for cancers with clear risk factors, like breast and colorectal Siegel et al. (2025), while cancers without early symptoms, such as pancreatic cancer, often progress silently and are detected late, with survival durations under one year (Blackford et al., 2024).
@@
We present CATCH-FM: CATch Cancer early with Healthcare Foundation Models, a cancer pre-screening tool that identifies high-risk patients using only their medical history. CATCH-FM is pretrained on large-scale longitudinal EHR data and finetuned on clinician-curated cancer cohorts. It directly operates on precise medical codes (ICD), learning general medical patterns through next-code prediction and finetuned to capture cancer risk signals reflected in patient medical history (Lee et al., 2021; 2022; Phan et al., 2020). Once trained, CATCH-FM can be deployed in EHR systems to predict cancer risk at low cost, supporting healthcare providers in deciding when and whom to screen.
@@
To facilitate the study of EHR foundation models in cancer pre-screening, we build NHIRD-Cancer, a cancer risk prediction benchmark, by sampling more than three million patients from the Taiwanese National Health Insurance Research Database (NHIRD) (Hsieh et al., 2019), a government de-identified and research-accessible EHR database. We allocate three million patients for pretraining, consisting of billions of medical events spanning two decades, and the remainder for finetuning and evaluation. We focus on three cancers based on their critical needs for early detection: pancreatic, liver, and lung cancers (Kim et al., 2024; Kukhareva et al., 2024; Thiele et al., 2024), and curate clinically reliable cancer cohorts after matching cancer patients with control-to-case groups.
@@
With billions of medical events available, we examine how compute budget (FLOPs), model size, and pretraining tokens affect performance. Our findings establish an EHR foundation model scaling law and confirm the benefit of large-scale pretraining on EHR data. Accordingly, we pretrain compute-optimal CATCH-FM models up to 2.4b parameters. When finetuned and evaluated for cancer risk prediction on NHIRD-Cancer, CATCH-FM consistently outperforms feature-based tree models and language models trained on the same data. Without loss of flexibility, it demonstrates strong predictive efficacy, achieving over 50% and 70% sensitivity in the first and subsequent target cancer cohorts at a 99% specificity cutoff, and reaching 50% and 80% AUPRC, respectively, offering strong reassurance when ruling out cancer (Grimes & Schulz, 2005).
@@
On the pancreatic cancer risk prediction task from Stanford Medicine EHRSHOT benchmark Wornow et al. (2023), CATCH-FM outperforms the prior state-of-the-art EHR foundation model, CLMBR Wornow et al. (2023), pretrained on millions of on-site patient records, while CATCH-FM is trained on data from drastically different populations, disease prevalence, and coding systems (ICD vs. SNOMED). Our analyses further demonstrate the robustness of CATCH-FM across different patient cohorts, preexisting conditions, and pre-screening configurations. Our interpretability analyses Gao et al. (2024) reveal that CATCH-FM identified not only known cancer risk factors but also non-trivial markers discovered in recent medical research.
@@ 2 Related Work
Cancer screening significantly improves prognosis and patient outcomes (Kim et al., 2024; Kukhareva et al., 2024; Thiele et al., 2024). Advancements in cancer treatment have made cancers more treatable and potentially curable, if they are detected in early stages (Chu et al., 2024; Hu et al., 2024; Kim et al., 2024; Liu et al., 2024). Recent advances in AI-powered medical imaging, like CT scans with multimodal models, have enhanced cancer screening accuracy, sometimes exceeding human-level sensitivity (Cao et al., 2023; Chen et al., 2023; Wang et al., 2024b; Xu et al., 2024a). The challenge is that medical imaging is resource-intensive, inaccessible for under-served populations, and not widely performable (Elmohr et al., 2024; Truhn et al., 2024; Vrudhula et al., 2024; Waite et al., 2021).
@@
Using electronic health records (EHR) to assess cancer risk is a promising path to improve cancer screening effectiveness and efficiency, i.e., by identifying patients with high cancer risk for healthcare professionals to make informed cancer screening decisions (Lee et al., 2021; 2022). Recent approaches have made attempts using feature-based machine learning models on large, task-specific datasets to detect cancer risks, and have shown the possibility of capturing cancer risk signals in EHR (Peduzzi et al., 2024; Placido et al., 2023b).
@@
Large language models pretrained on medical corpora are effective on text-oriented medical tasks (Clusmann et al., 2023) such as medical question answering (Singhal et al., 2025), clinical document summarization (Liu et al., 2025), and radiology report generation (Sun et al., 2024b). Recent research has also pretrained foundation models on structured EHR data (Choi et al., 2016; Gao et al., 2020; Yao et al., 2019), and explored their utilities in various clinical tasks (Savcisens et al., 2024; Sun et al., 2024a; Yang et al., 2023b). A challenge for EHR foundation models is that many EHR datasets only cover a slice of patients’ healthcare records (Faltys et al., 2021; Johnson et al., 2016; Pollard et al., 2018), e.g., MIMIC-IV mainly focuses on emergency departments and ICU encounters (Johnson et al., 2016). Foundation models pretrained on EHR slices are more effective in corresponding prediction tasks such as emergency department triage (Sun et al., 2024a) and ICU readmission risk (Jiang et al., 2023).
@@
Many tasks, like cancer risk prediction, require longitudinal EHR data to capture full patient history. Foresight uses two decades of data to pretrain Transformers for biomedical forecasting (Kraljevic et al., 2022). MOTOR trains a 143M-parameter model on time-to-event tasks using 2.7M patient records from 2014–2022, effectively predicting diagnosis time (Steinberg et al., 2023). CLMBR, a 141M-parameter Transformer, is pretrained on 2.57M Stanford Medicine records over three decades (Wornow et al., 2023), with 6.7K records released in the EHRSHOT benchmark for 15 few-shot prediction tasks (Wornow et al., 2023).
@@
Due to privacy and ethical constraints, large-scale longitudinal EHR data are hard to release publicly. Most EHR foundation models are pretrained on one or two hospital sites and are limited in scale to around 100M parameters (Steinberg et al., 2023; Wornow et al., 2023). Their advantage over feature-based models in cancer risk prediction remains unclear, for instance, while CLMBR performs well on procedure outcomes and lab predictions, it fails to outperform feature-based models in pancreatic cancer prediction on EHRSHOT (Wornow et al., 2023).
@@ 3.1 NHIRD Preliminary
Overview. The National Health Insurance Research Database (NHIRD) include electronic health record (EHR) of over 99.99% of Taiwan population. It includes decades of de-identified records from all healthcare encounters under the National Health Insurance program, with diagnoses, prescriptions, and procedures (Hsieh et al., 2019). Coding standards are applied uniformly across providers, ensuring consistency and correctness. A sample patient record is shown in Appendix Figure 6(b).
Availability. The sensitivity of longitudinal EHR data makes large-scale public access challenging (Guo et al., 2023). Among existing longitudinal EHRs (Kraljevic et al., 2022; Steinberg et al., 2023; Wornow et al., 2023), NHIRD is among the most accessible for open research. Eligible institutions in Taiwan and their international collaborators can access NHIRD with IRB approval or exemption under standardized protocols. Access is regulated by the Ministry of Health and Welfare to ensure strong privacy protections (Hsieh et al., 2019; Lin et al., 2018; Sung et al., 2020b). See Appendices A and B for details on de-identification, data quality, and access to NHIRD.
@@
| First Target Cancer Cohort | |||
| Total | 285,097 | 266,563 | 277,943 |
| Positive | 4,520 | 5,092 | 3,985 |
| Negative | 280,577 | 261,471 | 273,958 |
| Pos./Neg. Ratio | 1.61% | 1.95% | 1.45% |
| Subsequent Target Cancer Cohort | |||
| Total | 277,381 | 265,066 | 263,662 |
| Positive | 7,381 | 5,835 | 4,648 |
| Negative | 270,000 | 259,231 | 259,014 |
| Pos./Neg. Ratio | 2.73% | 2.25% | 1.79% |
@@ 3.2 NHIRD-Cancer Benchmark Curation
Data Statistics. We accessed a subset of NHIRD via collaboration with a Taiwanese medical school, with IRB exemption, as the study uses anonymized, non-human-subjects data. The dataset includes 3+ million patients from 1996–2013, covering 1.4 billion visits with diagnoses, treatments, and medications coded in ICD-9 (CDC, ). It provides 15+ years of physician-confirmed, government-validated history and over 8 billion medical codes. Summary statistics are in Table 1, with preprocessing and additional details in Appendix A. Our subset aligns with full NHIRD averages (18 vs. 16.5 visits per patient over 15 years) and is comparable in scale to the 3.67M-patient records from Stanford medicine used in CLMBR Wornow et al. (2023), which is not publicly available.
Targeted Cancers. We focus on pancreatic, liver, and lung cancers, which have high mortality rates and increased risk of metastatic cancer (De Visser & Joyce, 2023; Gupta & Massagué, 2006; Ji et al., 2023) where early detection offers substantial benefit (Kukhareva et al., 2024; Thiele et al., 2024; Blackford et al., 2024; Haue et al., 2024). The screening task is a binary classification: predicting whether a patient will later develop the target cancer. As suggested by clinicians, we focus on predicting target cancer one year after a clinical visit, enabling early intervention for proactive care. We set the available medical history per patient to be five years, following clinical standards. We explicitly split the benchmark into first target cancer (patients with no prior cancer history) and subsequent target cancer (patients with a history of other cancers but no prior target cancer) cohorts, and conduct experiments on models’ effacacy on both scenarios.
Case and Control Group. Guided by clinicians, we rigorously follow the standard case-control study methodology and previous cancer screening studies on NHIRD (Phan et al., 2020; Lee et al., 2021; 2022) to create case-control dataset. We use the first occurance of target cancer type as the positive case. The (negative) control group includes patients with no cancer diagnoses, ensuring a clear distinction between cases and controls, and are matched with the case group as follows.
@@
Our case-control matching goes beyond demographics by incorporating history length and timing, aligning patients more closely so differences reflect cancer status and enhancing validity.
NHIRD-Forward. We further collect an external dataset, referred to as NHIRD-Forward, from a different sample of the healthcare system at a distinct timeline from 2016-2021, 3 years after our NHIRD samples. The hospital only provided first liver and lung caner cohorts for testing. It follows the same NHIRD format while representing a different patient cohort with no overlap to NHIRD-Cancer, and is exclusively used to evaluate zero-shot generalization.
Dataset Processing. We use three-digit ICD-9 codes 157, 155 and 162 to define pancreatic, liver, and lung cancers, respectively. Control patients must not have cancer-related codes from 140–239 (top level of Neoplasms). To support broader cancer types and maximize early detection (Chen et al., 2020), we use top-level ICD codes to capture more cases. Table 2 lists benchmark statistics. All cancer diagnoses are confirmed by licensed physicians and validated by the government. More details in Appendix A.
@@ 4.1 Model
Patient Representation. In longitudinal structured EHR databases, the patient records are often represented by a coding system, for example, ICD CDC or SNOMED Cornet & de Keizer (2008), using a unique ID to encode each specific piece of information. A typical patient record can start with their demographic information, such as age and gender , followed by a sequence of chronologically ordered visits, . Each visit consists of medical events, , covering diagnoses, medications, procedures, and other clinical events.
@@
Architecture. CATCH-FM uses the standard decoder-only transformer architecture on top of the patient’s EHR token sequence . We use rotary positional embeddings (RoPE) (Su et al., 2023) to encode positional information for visits:
@@ 5 Experimental Methodologies
Dataset. We use the NHIRD-Cancer benchmark described in Sec. 3.2 for the pretraining, finetuning, and in-domain evaluation. We randomly partition the dataset, allocating 80% for pretraining and 20% for cancer screening benchmarks, ensuring no overlap between them. Each benchmark is further split into 80%/10%/10% for train/valid/test, using stratified sampling to maintain consistent class distributions for first and subsequent target cancer cohorts.
@@
To evaluate model generalization, We use NHIRD-Forward (Sec. 3.2) to assess zero-shot generalization where all patient cohort are used for evaluation. In addition, we use the EHRSHOT benchmark (Wornow et al., 2023), containing 6,739 Stanford Medicine patients (1990–2023), primarily coded in SNOMED, CPT, and RxNorm. It introduces significant distribution shifts from NHIRD: different healthcare systems, different medical codings, and different populations with different cancer prevalence. We use EHRSHOT’s official pancreatic cancer risk prediction task following their few-shot setting (1–128 positives). It is the only available cancer risk prediction task. We transfer EHRSHOT patients to NHIRD by exact code matching when possible and cosine-similarity embedding soft matching (threshold 0.98) for unmapped codes. Appendix C details the process.
@@
Evaluation metrics. NHIRD-Cancer evaluations use AUROC, AUPRC, specificity, and sensitivity, with the first two serving as the main metrics. AUROC measures the model’s ability to discriminate between patients with and without cancer across all decision thresholds. AUPRC captures the precision–recall trade-off in imbalanced tasks in cancer risk prediction. Specificity, the ratio of true negatives to all negatives, reflects reliability in identifying low-risk patients. Sensitivity, the ratio of true positives to all positives, reflects how well the model flags cancer-risk patients. When possible, we evaluate sensitivity at 99% specificity, a common (pre)screening threshold (Cao et al., 2023; Halner et al., 2023; Jopek et al., 2025), and a clinician-defined standard to reduce false positives, ensuring clinical trust.
@@
Baselines. We include standard tree-based models with bag-of-words as input (Bharati et al., 2023): XGBoost (Chen & Guestrin, 2016) and LightGBM (Ke et al., 2017), and well-established EHR Deep learning models (Wang et al., 2024a): StageNet (Gao et al., 2020) and RETAIN (Choi et al., 2016). We compare with pretrained language models, by converting medical codes into text and finetuning BioGPT (Luo et al., 2022) and Qwen2.5 (Yang et al., 2024) on the same data. For EHRSHOT, we compare CATCH-FM against methods reported in their leaderboard (Wornow et al., 2023).
@@
Implementation Details. We implement CATCH-FM using a standard architecture (Biderman et al., 2023), a decoder-only transformer with rotary position embeddings and flash attention, and pretrain from scratch on NHIRD. Medical codes, demographics, and special tokens form a 185,138 token vocabulary, with sequences capped at 2,048 tokens. Longer histories are chunked without overlap during pretraining and truncated for fine-tuning. More architecture, hyperparameters, and training details are provided in Appendix D.